OpenStudy (anonymous):

Can someone please explain for me. (2-3i)^2

OpenStudy (oaktree):

Sure. So we should write this out first, and we get\[(2-3i)(2-3i) = 4 - 6i - 6i +9i^2\]Following so far?

OpenStudy (oaktree):

@axelneel are you there?

OpenStudy (anonymous):

yes sorry i had another tab open. But that makes much more sense. why could wen not just use the rule (mn)^x=m^xn^x?

OpenStudy (anonymous):

it it because there is an addition symbol

OpenStudy (oaktree):

Yes. They are two different terms, not one.

OpenStudy (anonymous):

i see. what makes multiplying two terms okay to distribute the exponent to but not addition?

OpenStudy (oaktree):

Just think about it like they were two numbers. So\[(2*4)^2 = (8)^2 = 64\]But \[(2+4)^2 = 6^2 = 36\]They're totally different numbers on the inside.

OpenStudy (anonymous):

i see because of law of mutlication. 2^2*4^2=64

OpenStudy (oaktree):

Yeah.

OpenStudy (anonymous):

can you help me to simplfy -logx-logy+2logz into a single logorithm. where would i start?

OpenStudy (oaktree):

So we have\[-(\log x + \log y) + 2\log z = -(\log xy) + 2 \log z = \log z^2 - \log xy\]Can you continue from there? There's just one step left.

OpenStudy (oaktree):

Hint: \[\log a - \log b = \log \frac{ a }{ b }\]

OpenStudy (anonymous):

i seee, thank you very much

OpenStudy (anonymous):

would it be logz^2/logxy?